Your community newspaper, The Xpress, has fixed production costs of K70 per edition, and marginal printing and distribution costs of 40n/copy. The Xpress sells for 50n/copy.

a)   Write down the associated cost, revenue, and profit functions.                     

b)   What profit (or loss) results from the sale of 500 copies of The Xpress?         

c)    How many copies should be sold in order to break even?                      



1
Expert's answer
2021-07-21T09:44:28-0400

a)

Cost function , "C(x) = 40x + 70" K where x is number of newspaper.

Revenue function , "R(x) = 50x" K

Profit function, "P(x) = R(x) \& c(x) = 50x - 40x - 70"

So, "P(x) = 10x -70"


b).

When "x = 500"

Proft, "P(500) = 10(500) -70 = 5000 - 70 = 4930" K


(c) At break even "R(x) = C(x)"

So

"50x = 40x + 70\\
 50x -40x = 70\\
10x = 70\\
x = \frac{70}{10} =7"

For 7 copies to be sold to break even.